OpenStudy (anonymous):

0 = 250000(1.005)^n - 480000(1.005^n - 1) Solve for n

OpenStudy (lasttccasey):

0 = 250000(1.005)^n - 480000(1.005^n - 1) 0 = 480000 - 230000(1.005)^n <- Simplify. 480000/230000 = 1.005^n <- Get n by itself ln(480000/230000) = n*ln(1.005) <- Take Natural Log n = 147.5089 <- Solve for n